HUBBLE SPACE TELESCOPE DAILY REPORT #2120 PERIOD COVERED: 0000Z (UTC) 05/08/98 - 0000Z (UTC) 05/11/98 Daily Status Report as of 131/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Nine Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Three Sets of STIS/CCD 7803 (Daily Darks to Update Acquisition Bad Pixel Table Part III) The Space Telescope Imaging Spectrograph (CCD) was used to make daily dark calibration updates. These daily darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The observations were completed as planned, and no anomalies were reported. 1.3 Completed Two Sets of STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 7470 (A Snapshot Survey of Galactic Globular Clusters) The WF/PC-2 was used to make snap shot survey observations of the crowded center of the galactic globular cluster NGC-5634. The resulting color-magnitude diagram will include all stars down to about one magnitude below the main-sequence turnoff. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Seventeen Sets of NIC/2/3 7907 (NICMOS Pure Parallel Winter 97-98) The Near Infrared Camera (Cameras Two and Three) was used to make pure parallel spectrographic observations of sky regions when another instrument is prime. Just as the twelfth iteration of this proposal began, the NICMOS suspend/safe events described in 2.3 occurred, and this and the remaining iterations of this proposal did not execute. Otherwise, the observations were completed as planned, and no further anomalies were reported. 1.6 Completed Twelve Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.7 Completed NIC/1 7855 (The Evolution of Powerful Radio Galaxies) The Near Infrared Camera (Camera One, Two and Three) was used to make observations of the powerful radio galaxy PKS0941-08. Our proposed NICMOS observations combined with our multi-wavelength observing program will finally provide us with the data we need to build a picture of how powerful radio galaxies evolve. The observations were completed as planned, and no porblems were reported. 1.8 Completed Sixteen Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.9 Completed NIC/1/2/3 7822 (Grism Spectra of Centaurs and TNOs from 1-2 Microns) The Near Infrared Camera (Cameras One, Two and Three) was used to make grism spectra of object 199CQ29 from 1-2 microns. The surfaces of Centaurs and Trans-Neptunian Objects (TNOs) contain potentially unique records of early evolutionary processing {accretional and dynamical} in the outer solar system. The observations were executed as scheduled, and no anomalies were noted. 1.10 Completed WF/PC-2 7505 (The Hubble Diagram for Distant Supernovae Measuring Cosmic Deceleration and Global Curvature) The WF/PC-2 was used to make observations of Super Nova SN1998E in order to produce a precise Hubble diagram from measuring cosmic deceleration and global curvature. HST can make an essential contribution by providing photometry with substantially higher precision and better temporal coverage than can be obtained from the ground. The observations were completed as planned, and no anomalies were reported. 1.11 Completed Two Sets of NIC/2 7330 (The Fueling of Active Nucleii: A NICMOS Snapshot Survey) The Near Infrared Camera (Camera Two) was used to probe the dark matter distribution of NGC-3169 and NGC-3368 far from the center by observing intervening absorption from the MgII doublet (lambda 2800) in the spectrum of one background QSO located in a geometrically optimal position. HST observations are required because this feature is in a wavelength range not accessible from the ground, and the QSO magnitudes and expected strength of absorption make IUE observations unfeasible. The second iteration of this proposal did not execute due to the NICMOS suspend/safe events described in 2.3. Otherwise, the first iteration of the proposal completed as expected with no anomalies. 1.12 Completed Six Sets of S/C 7954 (FGS-1R Optimization) Desorption and gravity release effects have modified the morphology and modulation of the FGS-1R Transfer Function. This test will accomplish a re-optimization of the performance of FGS-1R by adjusting the FF3 mirror until the S curve at the center of the field of view is comparable to the optimized S curve obtained during SMOV. This test requires special data downlinks and uplinks for real-time commanding of the FF3. The test concluded with no reported problems. 1.13 Completed NIC/3 7892 (Narrow-band Imaging of BAL QSOs) The Near Infrared Camera (Camera Three) was used to make narrow band observations of the BL-Lac (BAL) Q1433-0025. The detection of extended emission-line gas structures in BAL QSOs would offer another means of investigating the BAL QSO phenomenon as well as establishing a new probe of material in galaxies at high redshift. The observations were completed as planned, and no problems were reported. 1.14 Completed Two Sets of WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.15 Completed Eight Sets of FGS 7684 (HST Astrometry Angular Resolution Test) The Fine Guidance Sensors (FGSs) were used to make astrometric observations of the binary star system ADS11300 as part of an astrometry angular resolution test. The ability of the HST FGS to detect and resolve close faint binary stars and measure their orbits is unique and unchallenged by current ground-based techniques. The observations were completed as planned, and no anomalies were reported. 1.16 Completed Twelve Sets of STIS/CCD 7911 (STIS Galactic Plane Pure Parallel Program) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude survey observations in parallel with the prime science observations. Short exposures with the 50CCD imaging mode will provide the wavelength zero-point calibration. These data will be placed immediately into the Hubble Data Archive. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 7431 (Snapshot Survey of Microlensed Source Stars) The WF/PC-2 was used to make snapshot survey images of microlensed source star 97-BLG-24. Gravitational microlensing surveys have revealed an excess of lensing events over theoretical predictions toward both the Large Magellanic Cloud and the Galactic Bulge. The observations were completed as planned, and no anomalies were reported. 1.18 Completed FGS 7955 (FGS/FGS Alignment After FGS1R Optimization) The fine Guidance Sensors were used to make observations of cluster NGC-5617 as part of an FGS/FGS alignment after FGS1R optimization. This test will acquire the data needed to calculate the FGS/FGS alignment calibration using an astrometric field. The observations were completed as planned, and no problems were reported. 1.19 Completed Five Sets of NIC/1 7867 (NICMOS Imaging of the Dusty CfA Seyfert Nuclei) The Near Infrared Camera (Camera One) was to be used to make observations of the Seyfert Galaxies NGC-5695 and NGC-5347. These data will allow us to determine if the Seyfert Type 1/2 dichotomy is best explained by the unified model, the evolutionary picture, or a combination of both of these scenarios. This proposal did not execute due to the NICMOS suspend/safe events described in 2.3. 1.20 Completed WF/PC-2 7277 (Calibration of the Cepheid P-L Relation with Observations of the Maser-Host Galaxy NGC) The WF/PC-2 was used to make observations of the maser-host galaxy NGC-4258 to study its Cepheid variable stars. These data will help refine the calibration of the Cepheid period-luminosity relationship. The observations were completed as scheduled, and no problems were reported. 1.21 Completed STIS/CCD/WF/PC-2/NIC/3 7388 (Black Holes and the Centers of Galaxies) The Space Telescope Imaging Spectrograph (CCD), the WF/PC-2, and the Near Infrared Camera (Camera Three) were used to continue our survey of centers of nearby galaxy NGC-3384. The major goal is to survey an unbiased set of galaxies with a potentially wide range of black hole masses. The sample has been carefully chosen to optimize leverage on the character of a proposed correlation of black hole mass with galaxy mass. This proposal was affected by the NICMOS suspend/safe events described in 2.3 (NIC/3 was utilized for a portion of the proposal). Otherwise, the observations were completed as scheduled with no further anomalies. 1.22 Completed FOC/96 6305 (Supernova Light Echoes II) The Faint Object Camera (F/96) was used to make observations of super nova SN1991T in order to study its light echoes. This is to carry out imaging polarization observations of candidate light echoes of historical supernovae to determine the distances to the host galaxies geometrically. The observations were completed as planned, and no anomalies were reported. 1.23 Completed WF/PC-2 7463 (Proper Motion of the Sagittarius Dwarf Galaxy) The unique spatial resolution of the WF/PC-2 was used to determine the proper motion of the Sagittarius dwarf galaxy. The observation was completed with no reported problems. 1.24 Completed NIC/1/2/3 7844 (H-2 Observations of the Galactic Center Circumnuclear Ring) The Near Infrared Camera (Cameras One, Two and Three) was to be used to make observations of the galactic center circumnuclear ring. These data will be used the distribution of molecular hydrogen in the circumnuclear ring that engulfs the inner two parsecs of the galaxy. This proposal did not execute due to the NICMOS suspend/safe events described in 2.3. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 32 Successful : 32 Scheduled Reacquisitions : 12 Successful : 12 2.2 FHST Updates: Scheduled : 65 Successful : 62 A roll update at 129/0955Z failed due to FHST #2. HSTAR 6466 was written, and subsequent guide star acquisitions were successful. Then, as described in HSTAR 6468, the full mode updates at 129/202706Z and 129/202951Z failed, both due to FHST #3. 2.3 Operations Notes: The FGS-1R fold flat #3 mirror was moved at 128/1130Z. After analysis, the final, contingency-only, mirror move for the FGS-1R Optimization was cancelled. The SSR correctable EDAC counts were cleared on five accosions using ROP SR-1. During one SSR playback of science data at 128/164529Z, there were 120 instances of unexpected, uncorrectable EDAC errors during an SAA contour. HSTAR 6463 was written. Per operations request, SSR fast memory monitors were set at 128/1932Z and an SSR1 interrupt log dump was performed at 129/0221Z. At 129/123403Z, the NICMOS instrument went into a supsended state due to the "NICMOS detector 2 VD voltage" ND2VDV, dropping below its lower limit. Additionally, NICMOS status buffer messages were issued. These messages and the limit violations were consistent with the signature for a TPG reset (which has been observed previously). HSTAR 6467 was written. This was followed about one minute later by the instrument going into the safe state at 129/123508Z as commanded by the NSSC-1 when the "NICMOS detector 1 DC offset voltage", ND1DCOFV, went out of limits low. HSTAR 6469 was produced. Analysis indicated that the detector 1 TPG did not reset correctly when the suspend was executed. Because the detector did not reset, the nominal negative voltage was still present, which then tripped the lower limit of zero for the boot mode. At 129/1940Z, flight operations personnel commenced the recovery procedures for NICMOS. The instument was brought to the boot state at 129/1945Z. At 129/1952Z, the instrument was returned to the operate state. The NICMOS detector 3 BB minus 15 voltage current flagged out of limits at 129/204404Z with a count of 1770 (limit is set to 1776, nominal to 1812). NICMOS went into another suspended state. During a telecon on May 10, it was decided that a possible explanation could be related to a combination of the MEB at a colder temperature due to the earlier safing event, and a possible degradation of the buffer box electronics due extended to radiation exposure. Another telecon is scheduled for today to discuss further analysis performed. A recommendation for recovery is expected to be made. The proposals described in 1.5, 1.11, 1.19, 1.21, and 1.24 were affected by the above events. The NSSC-1 status buffer was dumped at 130/2015Z per ROP NS-3. As described in HSTAR 6471, there was some engineering data loss (approximately 17 minutes) when the SSR was left misconfigured following a WF/PC re-transmission. HSTAR 6471 was written. Using ROP SR-3, the record pointer was reset. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH